The measure of the area of a rectangle is 4x2+7x-15. If hedemensions of the rectangle are representd by polynomials, find the demensions of the rectangle and then find its perimeter if each side is increased by 4 units.
Answer by stanbon(75887)

The measure of the area of a rectangle is 4x2+7x-15. If dimensions of the rectangle are representd by polynomials, find the dimensions of the rectangle and then find its perimeter if each side is increased by 4 units.
------------------------------
4x2+7x-15
Factor:
= 4x^2 +12x - 5x - 15
= 4x(x+3) -5(x+3)
= (x+3)(4x-5
These are the dimensions.
----------------------------
Increased dimentions: x+7 and 4x-1
Perimeter = 2(L + W)
= 2(x+7 + 4x-1)
= 2(5x+6)
= 10x+12
==========
Cheers,
Stan H.